The population of a small town at the beginning of the year was 7,000. Records show that during the year there were 5 births, 7 deaths, 28 new people moved into the town, and 12 moved out. what was the percent of increase or decrease in the town population?
My question is, how do you calculate percent increase or decrease?

Answer by richwmiller(17219)

If the town started out with 100 people and there was an increase of 12 people.
there would be and increase of 12/100 or 12%
So you first need to add and subtract all the changes to din d out if there are now more or fewer people.
+5-7+28-12=c
(c/7000)*100
14/7000*100=2*100/1000=2/10=.2 % increase
not 2 % but .2% which is 1/5 of 1 percent